OpenStudy (josedavid):

Write an equation of a line that passes through the point (2, 3) and is parallel to the line y = 3 over 2x + 5. y = negative 2 over 3x y = negative 2 over 3x − 7 y = 3 over 2x y = 3 over 2x + 7

OpenStudy (mathmale):

I explained earlier that your y = 3 over 2x + 5 is not appropriate as the equation of a straight line, and explaind that you must put the 3/2 in parentheses. Please type in the equation of this given straight line, doing it correctly this time. It must look like y = ( )x+b.

OpenStudy (josedavid):

is c

OpenStudy (josedavid):



OpenStudy (mathmale):

If y ou'd rather not cooperate, that's your decision. I 've asked you several times now to type in the given line properly. Would you do so, please? "I explained earlier that your y = 3 over 2x + 5 is not appropriate as the equation of a straight line, and explaind that you must put the 3/2 in parentheses. Please type in the equation of this given straight line, doing it correctly this time. It must look like y = ( )x+b."

OpenStudy (mathmale):

The given line is : _____________________________

OpenStudy (josedavid):

i dont get it

OpenStudy (mathmale):

Your "y = 3 over 2x + 5" is not correct. As explained before, you must show that the slope of this line is (3/2); it is not 3 over 2x. Please type out the given line.

OpenStudy (josedavid):

y = 3/2x + 5.

OpenStudy (mathmale):

Would you please enclose the slope 3/2 inside parentheses as I've asked you to do several times. y = ?

OpenStudy (josedavid):

example please

OpenStudy (mathmale):

"Your "y = 3 over 2x + 5" is not correct. As explained before, you must show that the slope of this line is (3/2); it is not 3 over 2x." Your slope here is a fraction, and so you'll need to enclose that fraction within parentheses to emphasize the fact that the slope is a fraction. y = ?

OpenStudy (mathmale):

Example: y = (1/3)x + 7. What is the slope?

OpenStudy (mathmale):

Your question presents the equation as \[y=\frac{ 3 }{ 2 }x+5.\]

OpenStudy (mathmale):

The slope is \[\frac{ 3 }{ 2 }\]

OpenStudy (josedavid):

ohhh

OpenStudy (mathmale):

If you're not going to use Equation Editor to present the slope properly, then enclose the '3/2" INSIDE PARENTHESES: \[y=\frac{ 3 }{ 2 }x+5~becomes~y=(3/2)x+5.\]

OpenStudy (josedavid):

ok

OpenStudy (mathmale):

What is the slope of this line? What is the slope of another line that is parallel to this line?

OpenStudy (josedavid):

y = negative 2 over 3x

OpenStudy (mathmale):

The answer to my question is: The slope of the given line is (3/2). Parallel lines have the same slope. So the slope of the new line you want is (3/2) also. This new line goes through the point (2,3). Use the point-slope form of the equation of a straight line to determine the equation of this new line. What does the point-slope formula look like?

OpenStudy (josedavid):

ok

OpenStudy (mathmale):

It's \[y-y _{1}=m(x-x _{1})\] Please substitute 3 for y1 and 2 for x1. What is the slope, m, of this new line? Remember, we've discussed this repeatedly and the slope, m, is given above.

OpenStudy (josedavid):

mmm

OpenStudy (mathmale):

On the left side of the equation above, you have \[y-y _{1}.\]Rewrite this by substituting 3 for y1. Please do that now.

OpenStudy (josedavid):

is hard

OpenStudy (mathmale):

If we do the substitution I asked for, y-y1 becomes y-3. Agree with that? This is the left side of your equation. Now, the right side is \[m(x-x _{1}).\]

OpenStudy (mathmale):

I ask you to substitute (3/2) (inside parentheses, please) for m and 2 for x1. Do this now, please.

OpenStudy (josedavid):

ok

OpenStudy (mathmale):

Good. Write it / type it here. The right side of your equation is m(x-x1). Replace m with (2/3) and x1 with 2, please. WRite it out now. type it here. share it with me.

OpenStudy (mathmale):

What do you now have for the right side of your equation? ( ? ) [x - ( ? ) ]

OpenStudy (mathmale):

jose david?

OpenStudy (josedavid):

yes is hard

OpenStudy (mathmale):

What do you need to know to be able to do what I'm asking? I asked you to replace m with (3/2) and replace x1 with 2. To help you along the way: after you replace m with (3/2), you get \[=(3/2) (x - [ ~~ ] )\]

OpenStudy (mathmale):

please type '2' inside the square brackets.

OpenStudy (josedavid):

2

OpenStudy (mathmale):

You need to type out the entire right side. The entire right side of your equation is = (3/2)(x-2). Have you any questions about this?

OpenStudy (josedavid):

yes what is the answer

OpenStudy (mathmale):

Can't give you the answer, Jose David. I will help you find your own answer. Your equation is now\[y-3= (3/2) (x-2).\]

OpenStudy (josedavid):

is c

OpenStudy (mathmale):

This is the answer, but it can be put into simpler form. All of the given possible answers are in simpler form, so you do have to put this into simpler form. Multiply the left and right sides of your equation by 2. You get: \[2(y-3)=3(x-2). \]

OpenStudy (josedavid):

how

OpenStudy (mathmale):

Multiplying this out, we get 2y-6=3x-6. Please simplify this. (cancel the 6's)

OpenStudy (josedavid):

ok

OpenStudy (mathmale):

I'm asking you to do something. "ok" lets me know you saw what I typed, but you haven't actually done the work I asked you to do. Simplify the above equation by subtracting 6 from both sides.

OpenStudy (josedavid):

i saw it

OpenStudy (mathmale):

subtract 6 from both sides: 2y-6=3x -6 Show me what's left.

OpenStudy (mathmale):

2y - 6 = 3x - 6 +6 = +6 -------------- Can you finish this, please? What is -6 + 6?

OpenStudy (josedavid):

12

OpenStudy (mathmale):

-6 + 6 is actually zero. +6 -6 ----- 0 Please finish this: 2y - 6 = 3x - 6 +6 = +6 --------------

OpenStudy (josedavid):

ohh

OpenStudy (mathmale):

Please finish that. -6 + 6 = 0. What does your final equation look like?

OpenStudy (josedavid):

umm

OpenStudy (mathmale):

On the left side: 2y - 6 + 6 --------- 2y + 0 2y

OpenStudy (mathmale):

any questions about that?

OpenStudy (josedavid):

no

OpenStudy (mathmale):

On the right side: 3x - 6 + 6 -------- Finish this, please.

OpenStudy (mathmale):

J-D: I need your input.